# Efficient Computation of Closed Substrings

### v2

@@ Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present a fast $\mathcal{O}(n\log n)$ time algorithm to compute all $\mathcal{O}(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. We also present a simple and space-efficient solution to compute all maximal closed substrings (MCSs) using the suffix array ($\mathsf{SA}$) and the longest common prefix ($\mathsf{LCP}$) array of $w[1..n]$. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
@@ 1 Introduction
The concept of open and closed strings was introduced by Fici in [12] to study the combinatorial properties of strings, with connections to Sturmian and Trapezoidal words. Then concepts of Longest Closed Factorization and Longest Closed Factor Array were introduced in [2], with improved solutions later proposed in [7]. Later, [3] introduced efficient algorithms for minimal and shortest closed factors. Related problems such as the $k$-closed string problem [1] and combinatorial problems on closed factors in Arnoux-Rauzy and $m$-bonacci words [14, 24] have also been explored.
It has been shown that a string contains at most $\mathcal{O}(n^{2})$ distinct closed factors [5]. More recently (in 2024), [23] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [20] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\mathcal{O}(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using weighted ancestor queries.
In [4] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [6] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later directly proved by Kosolobov in [18].
@@
Compute All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$, we compute all $\mathcal{O}(n^{2})$ closed substrings of $w$ in a compact representation using the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$), in $\mathcal{O}(n\log n)$ time.
Compute All MCSs: Given a string $w[1..n]$ we find all MCSs of $w$ using suffix arrays ($\mathsf{SA}$) and longest common prefix arrays ($\mathsf{LCP}$) in $\mathcal{O}(n\log n)$ time.
@@
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present algorithms to efficiently compute all closed substrings in a compact representation, as well as all MCSs, in $\mathcal{O}(n\log n)$ time. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. Finally, we present experimental analysis in Section 6.
@@ 2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n$. The empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if ji+1<nj-i+1<n. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover (for more details on covers, see the survey [19]).
A string $w[1..n]$ is periodic with period $t^{\prime}$ if $w[i]=w[i+t^{\prime}]$ holds for all $1\leq i\leq n-t^{\prime}$. The shortest period of $w$ is the smallest positive integer $t\leq t^{\prime}$ for which this condition holds. A periodic substring $w[i..j]$ of string $w[1..n]$ with shortest period $t$ is called a run if its length $j-i+1\geq 2t$ and its periodicity cannot be extended to the left or right, that is, $i=1$ or $w[i-1]\neq w[i+t-1]$, and $j=n$ or $w[j+1]\neq w[j-t+1]$.
@@ 3 Compact Representation of all Closed Substrings
A string $w[1..n]$ contains $\mathcal{O}(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\mathcal{O}(n^{2})$ closed substrings in Theorem 3.1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 2 in Section 4 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$ in a compact representation of size $\mathcal{O}(n\log n)$.
@@
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to j1j-1, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed. ∎
@@
Lemma 3(Theorem 1 in [18])
@@
Theorem 3.1 follows directly from Equation (1), Lemma 2 and 3.
Theorem 3.1
@@
An example of $\mathcal{C}(w)$ and the enumeration $\text{Closed}(w)$ for the string $w[1..11]=mississippi$ is given in Table 2 in the Appendix.
@@ 4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings in a compact representation and MCSs in $\mathcal{O}(n\log n)$ time. See Table 1 for an example of the $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$ in the Appendix.
Definition 1($\mathcal{MRC}$ Array [^2] This definition uses the notation introduced in Section 3.)
@@
We denote by $LCP_{max}$ the value stored in the root node of the AVL tree on the top of the stack. When a decrease in the $LCP$ value (or the end of the array) is encountered, the algorithm pops a collection of AVL trees from the stack, denoted as $\mathcal{L}_{suffixSets}$, such that all suffixes represented by an AVL tree key share the same longest common prefix equal to $LCP_{max}$. Since $LCP_{max}$ is the maximum $LCP$ value of all the AVL trees on the stack, popping this collection, $\mathcal{L}_{suffixSets}$, preserves the strictly decreasing order of ordered pairs in the $\mathcal{MRC}$ array. This collection is then merged into a single AVL tree using the Union operation presented in [16] [^3] The Union operation, defined in [16], builds upon foundational methods from [8, 9]., resulting in $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$. The Union operation also computes a list of ordered pairs representing maximal right-closed substrings via the $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation defined as follows:
The $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation then returns every pair of consecutive occurrences of substrings in the $\mathcal{L}_{suffixSets}$. These consecutive occurrences form the border of the maximal right-closed substring and are therefore added to the $\mathcal{MRC}$ array. Finally, the simple linear scan adds maximal right-closed substrings of length one. The algorithm clearly computes the maximal right-closed substrings in strictly decreasing order of length as required in the $\mathcal{MRC}$ array. Thus, the correctness of Algorithm 1 follows.
Algorithm 1 Compute $\mathcal{MRC}$ Array
Input: A string $w[1..n]$, its $\mathsf{SA}[1..n]$ and $\mathsf{LCP}[1..n]$.
Output: The $\mathcal{MRC}$ array of string $w[1..n]$.
[…]
2:while $stack\neq\emptyset\textbf{ or }i<n$ do
4:   $push(stack,AVLTree(\{\mathsf{SA}[i]\},\mathsf{LCP}[i]))$
5:   $i\leftarrow i+1$
6:  $\mathcal{L}_{suffixSets}\leftarrow\emptyset$
7:  $LCP_{max}\leftarrow LCP(top(stack))$
8:  while $stack\neq\emptyset\textbf{ and }LCP(top(stack))=LCP_{max}$ do
9:   $insert(\mathcal{L}_{suffixSets},pop(stack))$
10:  if $LCP_{max}\neq 0$ then
11:   $previousLCP\leftarrow LCP(top(stack))$
12:   $insert(\mathcal{L}_{suffixSets},pop(stack))$
13:   $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$ $\triangleright$ $\bigcup\mathcal{L}_{suffixSets}$ returns an AVL tree
14:   foreach $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ do $\triangleright$ Equation (5)
15:     $insert(\mathcal{MRC}[x],(y+LCP_{max}-x,LCP_{max}))$
16:   $LCP(\mathcal{T}_{suffixes})\leftarrow previousLCP$
17:   $push(stack,\mathcal{T}_{suffixes})$
18:for $i\leftarrow 1$ to $n$ do
19:  if $i=n\textbf{ or }(i<n\textbf{ and }w[i]\neq w[i+1])$ then
20:   $insert(\mathcal{MRC}[i],(1,0))$
21:return $\mathcal{MRC}$
Lemma 4([16])
Theorem 4.1
Algorithm 1 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$ in $\mathcal{O}(n\log n)$ time.
Proof
Algorithm to Compute $\mathcal{C}(w)$: Algorithm 2 computes the compact representation $\mathcal{C}(w)$, as in Equation (2), of all closed substrings of a string $w[1..n]$, by simply traversing all $n$ lists in the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). Since the $\mathcal{MRC}$ array is of size $\mathcal{O}(n\log n)$ we get Theorem 4.2.
Algorithm 2 Compute $\mathcal{C}(w)$
Input: $\mathcal{MRC}$ array of strings of $w[1..n]$
Output: Compact representation $\mathcal{C}(w)$ as in Equation (3)
[…]
2:for $i\leftarrow 1$ to $n$ do
3:  for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
5:   if $j=1$ then
6:     $\ell p_{j}\leftarrow 1$ $\triangleright$ Equation (2)
7:   else
8:     $(\ell r_{j-1},\ell b_{j-1})\leftarrow\mathcal{MRC}[i][j-1]$
[…]
11:return $\mathcal{C}$
Theorem 4.2
Algorithm to Compute All MCSs: Algorithm 3 computes all MCSs of a string $w[1..n]$ by iterating through the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). For each index $i$ in $w$ (where $1\leq i\leq n$), the algorithm examines each maximal right-closed substring represented in $\mathcal{MRC}[i]$ and verifies whether the substring is also maximal left-closed using a constant time check $i=1$ or $(i>1\text{ and }w[i-1]\neq w[i+|r|-|b|-1])$. If the condition is satisfied, the substring is added to the MCS list. Since scanning through the $\mathcal{MRC}$ array takes $\mathcal{O}(n\log n)$ time, and each check is performed in constant time, the algorithm computes all MCSs in $\mathcal{O}(n\log n)$ time, establishing Theorem 4.3.
Algorithm 3 Compute all MCSs
Input: A string $w[1..n]$ and $\mathcal{MRC}$ array
Output: All MCSs
[…]
2:for $i\leftarrow 1$ to $n$ do
3:  for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
5:   if $i=1$ or $(i>1\textbf{ and }w[i-1]\neq w[i+\ell r-\ell b-1])$ then
6:     $insert(mcsList,w[i..i+\ell r-1])$
7:return mcsListmcsList
Theorem 4.3
5 MCSs in a Fibonacci Word
Bucci et al. [10] study the open and closed prefixes of Fibonacci words and show that the sequence of open and closed prefixes of a Fibonacci word follows the Fibonacci sequence. De Luca et al. [11] explore the sequence of open and closed prefixes of a Sturmian word. In [14], the authors define and find the closed Ziv–Lempel factorization and classify closed prefixes of infinite $m$-bonacci words. In this section, we extend previous work on closed substrings in Fibonacci words by deriving a formula to compute the number of MCSs they contain.
The non-singleton MCSs are either runs or maximal gapped repeats in which the arm occurs exactly twice [4]. For a Fibonacci word $f_{n}$, let $SM(f_{n})$, $R(f_{n})$, and $GM(f_{n})$ denote the number of singleton MCSs, runs, and gapped-MCSs, respectively. Therefore, the total number of MCSs in $f_{n}$ is given by $M(f_{n})=SM(f_{n})+R(f_{n})+GM(f_{n})$.
Lemma 5
The no. of singleton MCSs in a Fibonacci word $f_{n}$, for $n\geq 4$, is:
Proof
Lemma 6(Theorem 2 in [17])
To compute $GM(f_{n})$, we count the number of maximal gapped repeats whose arms occur exactly twice, as these are, by definition gapped-MCSs. In Theorem 5.1 we first show that the only gapped-MCSs in a Fibonacci word are occurrences of the substring $101$ that are also maximal gapped repeats. Then, we count all such occurrences in $f_{n}$.
Lemma 7(Theorem 1 in [27])
Lemma 8
Proof
Lemma 9(Theorem 2 in [27])
Lemma 10(Lemma 8 in [27])
Lemma 11
For $n\geq 5$, the border $p_{n-1}$ of $p_{n}$ is the longest proper cover of $p_{n}$.
Proof
Lemma 12(Lemma 2 in [21])
Lemma 13
Proof
Theorem 5.1
Proof
Lemma 14
The no. of gapped-MCSs in a Fibonacci word $f_{n}$, for $n\geq 5$, is:
Proof
We prove the Lemma by induction on $n$. By direct counting for the base cases, we verify that $GM(f_{5})=1$ and $GM(f_{6})=2$. For the inductive step, assume that Equation (7) holds for $k\geq 7$. For $n\geq 3$, $f_{3}=101$ is a prefix of every $f_{n}$, and by Theorem 5.1 every gapped-MCS in a Fibonacci word $f_{n}$, for $n\geq 5$, corresponds to an occurrence of the substring $101$ that is also a maximal gapped repeat. In $f_{k+1}=f_{k}f_{k-1}$, any new occurrences of $101$ in $f_{k+1}$ can only occur beginning in $f_{k}$ and ending in $f_{k-1}$. Moreover, an occurrence of $101$ as a suffix of $f_{k}$ or a prefix of $f_{k-1}$ may be lost by extension. Below we examine all such cases.
Theorem 5.2
Proof
6 Experimental Analysis
All experiments were carried out on a server with an Intel Xeon Gold 6426Y CPU (64 cores, 128 threads, 75 MiB L3 cache), 250 GiB RAM, running RHEL 9.5 (kernel version 5.14.0-503.26.1). The code was implemented in C++ and compiled using GCC 11.5.0. The implementations for computing the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$) were sourced from the libsais library [13], which is based on contributions from [15, 22, 25, 26]. The implementation is available at https://github.com/neerjamhaskar/closedStrings.
For $\sigma=2,3$ and $4$, we generate all strings of length $n$ for $1\leq n\leq 20$ and compute their MCSs. The results are plotted in Figure 2 in the Appendix. As observed, the maximum number of MCSs increases with the size of the alphabet $\sigma$ for strings of the same length. In total, the number of strings analyzed is equal to $\sum_{i=1}^{20}(2^{i}+3^{i}+4^{i})\approx 1.47$ trillion. These exhaustive computations raise a natural question: Given $n$ and $\sigma$, can we algorithmically construct a string that has the maximum possible number of MCSs?
{credits}
6.0.1 Acknowledgements
We thank Simon J. Puglisi for introducing us to the MCS problem using $\mathsf{SA}$ and $\mathsf{LCP}$, which led to Theorem 4.3. We are grateful to the reviewers for their valuable feedback.
Appendix

### v3

# Efficient Algorithms to Compute Closed Substrings

@@ Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present fast $\mathcal{O}(n\log n)$ time algorithms to compute all $\mathcal{O}(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. These simple and space-efficient algorithms also compute maximal closed strings. Furthermore, we compare the performance of these algorithms and identify classes of strings where each performs best. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
keywords:
Closed Strings, Maximal Closed Substrings, Fibonacci Words
@@ 1 Introduction
The concept of open and closed strings was introduced by Fici in [fici_2011_intro] to study the combinatorial properties of strings, with connections to Sturmian and Trapezoidal words. Then concepts of Longest Closed Factorization and Longest Closed Factor Array were introduced in [Badkobeh2014ClosedFactorization], with improved solutions later proposed in [Bannai2015SPIRE]. Later, [Badkobeh2016DAM] introduced efficient algorithms for minimal and shortest closed factors. Related problems such as the $k$-closed string problem [Alamro2017EANN] and combinatorial problems on closed factors in Arnoux-Rauzy and $m$-bonacci words [JAHANNIA202232, PARSHINA201922] have also been explored.
It has been shown that a string contains at most $\mathcal{O}(n^{2})$ distinct closed factors [Badkobeh2015LATA]. More recently (in 2024), [PARSHINA2024114315] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [Mieno2025SOFSEM] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\mathcal{O}(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using weighted ancestor queries.
In [Badkobeh_2022_SPIRE] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [BADKOBEH2026115628] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later directly proved by Kosolobov in [kosolobov2024closedrepeats].
@@
Compute All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$, we compute all $\mathcal{O}(n^{2})$ closed substrings of $w$ in a compact representation in $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ auxiliary space.
Compute All MCSs: Given a string $w[1..n]$ we find all MCSs of $w$ in $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ auxiliary space.
@@
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present two algorithms to compute all closed substrings, including all MCSs. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. We then present experimental analysis of the proposed algorithms in Section 6. Finally, we conclude with future work and open problems in Section 7.
An earlier and abridged form of this work appeared in SPIRE 2025 [KJAIN_SPIRE_2025]. In this extended version, we present an alternate algorithm to compute all closed substrings, including all MCSs, analyze the space complexities of both algorithms, and provide extensive experimental results comparing their performance across different classes of strings.
@@ 2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n$. The empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if ji+1<nj-i+1<n. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover (for more details on covers, see the survey [NEERJA_STRING_COVERING]).
A substring $w[i..j]$ is primitive if it is not of the form $u^{e}$ for some $u$ with $e\geq 2$. The integer $e$ is the exponent of $w$. A substring $w[i..j]$ is a repetition in $w$ if it is of the form $u^{e}$ for some primitive $u$ with $e\geq 2$. A repetition $w[i..i+et-1]=u^{e}$ is a maximal repetition if $i=1$ or $w[i-1..i+et-2]\neq u^{e}$, and i+et1=ni+et-1=n or $w[i+1..i+et]\neq u^{e}$. A string $w[1..n]$ is periodic with period $t$ if it can be written as $w=u^{e}u^{\prime}$, where $u$ is primitive, $t=|u|$, $u^{\prime}$ is a (possibly empty) prefix of $u$, and either $e\geq 2$ or $e=1$ with $u^{\prime}$ non-empty. A run is a periodic substring $w[i..j]=u^{e}u^{\prime}$ of $w[1..n]$, where $e\geq 2$ and its periodicity cannot be extended to the left or to the right; that is, $i=1$ or $w[i-1]\neq w[i+t-1]$, and $j=n$ or $w[j+1]\neq w[j+1-t]$.
@@
Overview of Crochemore’s Maximal Repetition Algorithm
We provide a brief overview of Crochemore’s well-known algorithm for computing maximal repetitions [CROCHEMORE1981244]–which we refer to as CMR Algorithm. This algorithm serves as the basis for our Algorithm 2.
For a given string $w[1..n]$, we consider the positions within the string to be $i,j\in\{1,2,\ldots,n\}$. Let $1\leq k\leq n$ denote the length of a substring, also referred to as a level in the CMR Algorithm. Formally, we define an equivalence relation $\approx_{k}$ on the set of positions $\{1,2,\ldots,n\}$ by
$i\approx_{k}j\Longleftrightarrow w[i..i+k-1]=w[j..j+k-1].$
The corresponding equivalence classes (maintained as ordered-lists) of $\approx_{k}$ form a partition of the set of positions $\{1,2,\ldots,n\}$. For each position $i$, we denote by
$i]_{k}=\{\,j\mid j\approx_{k}i\,\}$
the equivalence class of $i$ under $\approx_{k}$. Therefore, each class $i]_{k}$ consists of the starting positions of all occurrences of a substring of length $k$ in $w$. Let $U_{k}$ be the set of all distinct substrings of $w$ of length $k$. Then, the set of equivalences classes at each level $k$ is defined as
$C_{k}=\Bigl\{\{j\mid w[j..j+k-1]=u\}\Bigm|u\in U_{k}\Bigr\}.$
Figure 3 shows the classes of $\approx_{k}$ where $1\leq k\leq 5$ for the string $w[1..11]=mississippi$.
The equivalence classes at level $k=1$ can be computed by a simple scan of the string from left to right, grouping together all positions that correspond to the same symbol in $\Sigma$.
The computation of the equivalence classes at level $k+1$ from the classes at the previous level, $k$, is called the refinement process, which is illustrated with arrows in Figure 3. Once the class size at any level becomes $1$ it cannot be refined further. When all classes contain just one element (singleton classes) the algorithm terminates.
To compute the set of classes at level $k+1$ from the classes at level $k$, we refine the classes at level $k$ one by one. Let $C^{\prime},C^{\prime\prime}\in C_{k}$ such that $C^{\prime}\neq C^{\prime\prime}$. Then, to refine a class $C^{\prime}$, we consider indices $i,j\in C^{\prime}$ such that $i\neq j$. If $i+1,j+1\in C^{\prime\prime}$, then $i$ and $j$ remain in the same class; otherwise they are moved to different classes. Once we refine all classes at level $k$, we get all the classes at level $k+1$.
The $\mathcal{O}(n\log n)$ time complexity, in Lemma 1, lies in the choice of classes for refinement. For each family of classes (classes that were formed as a refinement of a class on the previous level), we identify the largest class and call all the others small. By using only small classes for refinement to get all the equivalence classes in the next level, $\mathcal{O}(n\log n)$ complexity is achieved, as each element belongs only to $\mathcal{O}(\log n)$ small classes.
Lemma 1 (Theorem 7 in [CROCHEMORE1981244]).
The CMR Algorithm correctly computes all the maximal repetitions in a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
@@ 3 Compact Representation of all Closed Substrings
A string $w[1..n]$ contains $\mathcal{O}(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\mathcal{O}(n^{2})$ closed substrings in Theorem 1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 3 in Section 4 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$ in a compact representation of size $\mathcal{O}(n\log n)$.
@@
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to j1j-1, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed.
@@
Lemma 4 (Theorem 1 in [kosolobov2024closedrepeats]).
@@
Theorem 1 follows directly from Equation (1), Lemma 3 and 4.
Theorem 1.
@@
An example of $\mathcal{C}(w)$ and the enumeration $\text{Closed}(w)$ for the string $w[1..11]=mississippi$ is given in Table 2.
@@ 4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings in a compact representation and MCSs in $\mathcal{O}(n\log n)$ time. See Table 1 for an example of the $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$. We present two algorithms: Algorithm 1 based on $\mathsf{SA}$ & $\mathsf{LCP}$ array and Algorithm 2 based on Crochmore’s equivalence classes approach.
Definition 1 ($\mathcal{MRC}$ Array [^2] This definition uses the notation introduced in Section 3.).